OpenStudy (anonymous):

FAN AND MEDAL*** The table below shows some values of f(x) and g(x) for different values of x:

OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x: x f(x) = 2x + 24 g(x) = 5(0.5)x −2 20 −1 0 5 1 2.5 2 28

OpenStudy (anonymous):

x f(x) = 2x + 24 g(x) = 5(0.5)x −2 20 −1 0 5 1 2.5 2 28

OpenStudy (anonymous):

Complete the chart and determine the solution of the equation f(x) = g(x). A)x = −2 B)x = −1 C)x = 1 D)x = 20

OpenStudy (anonymous):

@saseal @DanJS @DaBest21 can you guys help me

OpenStudy (anonymous):

plug in the numbers to the function and calculate the output

OpenStudy (danjs):

yes as saseal says, just use the x value in each function to calculate the missing pieces of the chart... f(x) = g(x) can then just be read off the chart

OpenStudy (anonymous):

okay, let me see

OpenStudy (danjs):

g(x) = 5*(0.05)^x right, not multiplied by x

OpenStudy (anonymous):

x f(x) = 2x + 24 g(x) = 5(0.5)x -2 20 ? -1 ? ? 0 ? 5 1 ? 2.5 2 28 ?

OpenStudy (anonymous):

its the answer : x=-2?

OpenStudy (danjs):

g(-2) = 5*(0.5)^(-2) = 20

OpenStudy (anonymous):

f(x) = 2x + 24 use x = 2 and plug in 2 for every x, f(2) = 2(2) + 24 f(2) = 4 + 24 = 28 f(2) = 28 g(x) = 5(0.5)^x g(2) = 5(0.5)^2 g(2) = 5(0.25) g(2) = 1.25

OpenStudy (anonymous):

f(-2) = 2(-2) + 24 f(-2) = 20 g(-2) = 5*4 g(-2) = 20 Its x = -2

OpenStudy (danjs):

yep, looks good

OpenStudy (danjs):

why you need help.. hah you got it all

OpenStudy (anonymous):

i just want to make sure about my answer but thank you!!!

OpenStudy (danjs):

welcome

OpenStudy (anonymous):

remember to put this sign ^ for power